# Exploiting Completeness Perception with Diffusion Transformer for Unified 3D MRI Synthesis

### v1

@@ Abstract
Missing data problems, such as missing modalities in multi-modal brain MRI and missing slices in cardiac MRI, pose significant challenges in clinical practice. Existing methods rely on external guidance to supply detailed missing state for instructing generative models to synthesize missing MRIs. However, manual indicators are not always available or reliable in real-world scenarios due to the unpredictable nature of clinical environments. Moreover, these explicit masks are not informative enough to provide guidance for improving semantic consistency. In this work, we argue that generative models should infer and recognize missing states in a self-perceptive manner, enabling them to better capture subtle anatomical and pathological variations. Towards this goal, we propose CoPeDiT, a general-purpose latent diffusion model equipped with completeness perception for unified synthesis of 3D MRIs. Specifically, we incorporate dedicated pretext tasks into our tokenizer, CoPeVAE, empowering it to learn completeness-aware discriminative prompts, and design MDiT3D, a specialized diffusion transformer architecture for 3D MRI synthesis, that effectively uses the learned prompts as guidance to enhance semantic consistency in 3D space. Comprehensive evaluations on three large-scale MRI datasets demonstrate that CoPeDiT significantly outperforms state-of-the-art methods, achieving superior robustness, generalizability, and flexibility. The code is available at https://github.com/JK-Liu7/CoPeDiT.
@@ 1 Introduction
Magnetic resonance imaging (MRI) is widely used in clinical medicine due to its non-invasive nature and ability to visualize detailed tissue properties [8]. For instance, multi-modal brain MRIs provide complementary information regarding brain anatomy and pathology [27], while volumetric cardiac MRIs reveal detailed cardiac structure and tissue characteristics [7]. However, brain and cardiac MRIs often suffer from missing data in real-world clinical settings [50]. For brain MRIs, certain modalities may be unavailable, while for cardiac MRIs, some slices may be missing, which are commonly caused by limited scanning time, image corruption, and variations in acquisition protocols [33].
Figure 1: The motivation of CoPeDiT. (a) Comparison of completeness perception between CoPeDiT and prior methods. (b) Prompt tokens offer more effective guidance than binary mask codes. (c) CoPeVAE enables more discriminative latent representations. (d) Prompt tokens yield more semantically consistent attention maps in MDiT3D blocks, highlighting stronger interactions between similar modalities (e.g., T1-T1ce, T2-FLAIR).
To address this, generative models have been developed to infer missing data from observed inputs [19, 11]. Existing paradigms rely on auxiliary embeddings, e.g., binary mask codes, as prior knowledge to encode missing patterns (e.g., severity, type, and position) [24, 14, 21, 6, 29]. Nonetheless, this approach has several inherent limitations that hinder its practical applicability. First, in real-world deployments, input images may originate from different hospitals, scanners, or time points. As missing patterns are intrinsically unpredictable, providing manually defined masks in advance is impractical [49, 41]. Second, this over-reliance on predefined masks causes generative models to be unaware of the missing state and unable to distinguish between different modalities or spatial characteristics, thereby reducing robustness to variable and unseen incomplete patterns [22, 20, 51]. Models conditioned on fixed mask representations often suffer from limited generalizability, exhibiting noticeable drop in synthesis quality when encountering unfamiliar missing sites [13, 2, 32]. Finally, binary mask codes lack semantic richness and adaptability, leading to a rigid guidance that weakens spatial alignment and semantic coherence in the synthesized output [37, 17, 44].
@@
Intuitively, generative models should be capable of inferring and detecting the incomplete state spontaneously, rather than relying on externally manual guidance [17, 12]. Motivated by this, we pose a central question: ‘Can we empower the model with the ability to perceive missing states on its own?’ In light of this, we exploit a long neglected and underexplored property of generative models in medical imaging, i.e., completeness perception, to enhance flexibility and generalizability under arbitrary missing MRI conditions, as illustrated in Fig. 1 (a). Our fundamental insight is to enable the generative model to recognize the fine-grained incomplete state information in a self-perceptive manner, and to leverage this understanding as internal prompts to guide the generation process. We hypothesize that, for diffusion models, such self-guided prompts may serve as an effective alternative to manually defined masks, and potentially offer even stronger guidance signals (Fig. 1 (b)). The main reason lies in the fact that this self-perceptive strategy encourages the model to learn both global and local anatomical structures and lesion patterns at coarse and fine levels, thus enabling more semantically coherent generation of the missing MRI regions during synthesis (Fig. 1 (c)(d)) [23, 18].
@@
Driven by our motivation, we propose CoPeDiT, a 3D latent diffusion model (LDM) framework for unified 3D MRI synthesis. Technically, our framework builds on two core components: (i) Unlike prior approaches that require explicit missing indicators, a novel tokenizer with a completeness perception function, CoPeVAE, is proposed to autonomously assess the integrity of modalities or volumes through tailored self-supervised pretext tasks. By detecting anatomical structures and variations in lesion patterns, CoPeVAE develops a comprehensive understanding of 3D MRIs. This enables CoPeDiT to eliminate the need for manual intervention with flexible adaptability, enhancing the method’s autonomy and improves its feasibility for real-world clinical deployment with diverse missing patterns. (ii) A novel diffusion transformer architecture, MDiT3D, is developed to unlock the potential of Diffusion Transformers (DiT) [34] for medical image synthesis, offering a new paradigm for generating high-quality volumetric data. MDiT3D leverages tokenization and attention mechanisms to model long-range and irregular contextual dependencies in high-dimensional, structurally complex 3D MRIs. It supports conditional generation guided by completeness-aware prompts, which provide semantic-meaningful guidance on the synthesis process. The learned prompt tokens comprehensively absorb information from available modalities or slices during the diffusion process and guide the generation of the missing ones. Incorporating the above two innovations, our architecture not only enables adaptive self-guidance synthesis in a mask-free manner which exhibits strong robustness and generalizability, but also demonstrates improved structural coherence and enhanced preservation of fine-grained anatomical details.
Our main contributions are summarized as follows:
@@
  • We propose a unified framework, dubbed CoPeDiT, for both 3D brain and cardiac missing MRI synthesis under arbitrary incomplete scenarios, without the need for explicit external indicators as guidance.
@@
  • We present MDiT3D, a novel latent diffusion transformer for 3D MRI synthesis, featuring a dedicated architecture for capturing modality/spatial relationships and leveraging generated prompts as conditional guidance.
  • Extensive experiments on three datasets demonstrate that our model surpasses state-of-the-art (SOTA) methods, achieving generalizability, robustness, and real-world clinical applicability.
@@ 2 Related Work
MRI Synthesis. Recent efforts focus on unified MRI synthesis using generative models such as Generative Adversarial Networks (GAN) and diffusion models [52, 56, 1, 36]. For example, MMGAN [43] employs a single network to handle diverse missing configurations, while Hyper-GAE [53] introduces a shared hyper-encoder and decoder architecture with a graph-based fusion module for reconstructing missing brain MRIs. MMT [24] proposes to leverage Swin Transformers [25] to capture inter-modal dependencies. M2DN [29] presents a diffusion-based method that jointly models all modalities using a multi-input multi-output framework. Despite these advances, most existing methods still rely on externally provided masks to encode missing patterns in randomly incomplete cases. In contrast, CoPeDiT explores self-perceptive capability of generative models to autonomously recognize data completeness, enabling more flexible and high-fidelity MRI synthesis.
Diffusion Models. Diffusion models have shown strong generative capacities across various computer vision tasks [26, 28, 55]. Denoising Diffusion Probabilistic Models [16] generate images by gradually adding Gaussian noise and learning to reverse the process. LDM [39] improve efficiency by operating in a compressed latent space using autoencoders. Recently, Diffusion transformers [34] have been effective alternatives to U-Net [40] in LDMs, achieving competitive performance in natural image generation. However, most prior medical image generation approaches still rely on U-Net-based architectures as the diffusion backbone, leaving the potential of transformer-based diffusion models underexplored in this domain [48, 31]. In this paper, we devise MDiT3D, by replacing the U-Net backbone with a diffusion transformer and incorporating task-specific architectural modifications, MDiT3D is capable of generalizing to 3D MRI data and capturing long-range dependencies of high-dimensional 3D MRIs.
Figure 3: The overview framework of MDiT3D. PE and RoPE denote positional embeddings and rotary position embeddings [45, 54], respectively.
3 Methodology
@@ 3.1 Notations
We consider two 3D MRI synthesis tasks: brain missing modality synthesis and cardiac missing slice synthesis. To unify the formulation, let $\mathcal{M}=\{\mathbf{x}^{i}\}_{i=1}^{m}$ denote a complete MRI sample, where $m$ is the total number of modalities or slices. The available and missing subsets are denoted as $\mathcal{M}^{S}=\{\mathbf{x}^{s_{i}}\}_{i=1}^{s}$ and $\mathcal{M}^{C}=\{\mathbf{x}^{c_{i}}\}_{i=1}^{c}$, respectively, with $m=s+c$. (1) Brain missing modality synthesis: $\mathcal{M}$ represents all brain MRIs with $m$ modalities, and the task is to infer $\mathcal{M}^{C}$ from $\mathcal{M}^{S}$. (2) Cardiac missing slice synthesis: $\mathcal{M}$ corresponds to a volume with $m$ slices, where missing slices $\mathcal{M}^{C}$ are consecutive, non-overlapping with $\mathcal{M}^{S}$, and are to be reconstructed from the available ones. Notably, our experimental setup considers randomly generated incomplete cases from all the possible combinations as input, with a varying number and length of missing modalities or slices, aiming to mimic real-world clinical environments.
@@ 3.2 Stage I: Completeness Perception Tokenizer
The core idea of CoPeVAE (Fig. 2) is that detecting the completeness of high-resolution MRI data enforces the model to perceive both global anatomy and local lesion patterns, thereby producing high-quality prompts as diffusion guidance. Building upon VQGAN [47, 10], we deploy a 3D autoencoder jointly trained with self-supervised pretext tasks. Each task employs a prompt encoder, denoted as $\mathcal{F}_{1},\mathcal{F}_{2}$ and $\mathcal{F}_{3}$, to transform latent tokens (learned by the encoder $\mathcal{E}$) into low-dimensional prompt tokens. All prompt encoders contain 3D Conv layers followed by spatial average pooling. Afterwards, task-specific projection heads are applied for multi-granular classification and contrastive learning.
@@
Task 1. Missing Number/Length Detection. Equipped with global contextual awareness, the tokenizer is capable of answering the question that how many modalities/slices are missing in the incomplete input. This task aims to enable the tokenizer to identify the severity of incompleteness by perceiving the global context of MRIs, thereby learning modality/spatial attributes in a coarse-grained manner. We formulate this task as an (m1)(m-1)-class classification task and define the loss using the cross-entropy as
$\small\mathcal{L}_{d}=\mathcal{L}_{\text{cls}}(\mathcal{H}_{1}(\mathcal{F}_{1}(\mathbf{z}^{s}))_{c}).$ (1)
where $\mathcal{F}_{1}$ and $\mathcal{H}_{1}$ represent the prompt encoder and the projection head for missing detection, respectively. The learned prompt tokens $\mathbf{p}^{d}=\mathcal{F}_{1}(\mathbf{z}^{s})$ are rich in information about the severity of missing state.
@@
$\small\mathcal{L}_{p}=\mathcal{L}_{\text{cls}}(\mathcal{H}_{2}(\mathcal{F}_{2}(\mathbf{z}^{s}))_{I}).$ (2)
@@
Task 3. Missing Modality/Slice Assessment. Motivated by the observation that modalities or slices from the same scan share more similar anatomical and textural context than those from different scans, we adopt an inter-modal/inter-slice contrastive learning scheme [38] to serve as a missing data estimator. Specifically, we take the incomplete latent tokens $\mathbf{z}^{s}$ as anchor, the corresponding missing latent $\mathbf{z}^{c}$ from the same subject as positives, and latent tokens $\mathbf{z}^{c}_{-}$ from different subjects as negatives. The contrastive loss is defined as follows:
$\mathcal{L}_{c}=-\log\frac{\varphi\left(\mathcal{H}_{3}(\mathbf{p}^{s}),\mathcal{H}_{3}(\mathbf{p}^{c})\right)}{\varphi\left(\mathcal{H}_{3}(\mathbf{p}^{s}),\mathcal{H}_{3}(\mathbf{p}^{c})\right)+\sum\varphi\left(\mathcal{H}_{3}(\mathbf{p}^{s}),\mathcal{H}_{3}(\mathbf{p}^{c}_{-})\right)}.$
where $\mathbf{p}^{s}=\mathcal{F}_{3}(\mathbf{z}^{s})$, $\quad\mathbf{p}^{c}=\mathcal{F}_{3}(\mathbf{z}^{c})$, $\varphi(a,b)=\exp(a\cdot b/\tau)$, and $\tau$ is the temperature parameter. This contrastive learning scheme encourages the model to focus on inter-modal/slice contextual differences, thus improving anatomical coherence and fine-grained detail preservation. The overall loss of CoPeVAE is formulated as
$\small\mathcal{L}_{\text{tok}}=\mathcal{L}_{\text{rec}}+\lambda(\mathcal{L}_{d}+\mathcal{L}_{p}+\mathcal{L}_{c}).$ (4)
where $\lambda$ is the weighting coefficient, and $\mathcal{L}_{\text{rec}}$ denotes the reconstruction loss that consists of a pixel-wise L1 loss, vector quantization loss, adversarial loss, and perceptual loss.
Table 1: Quantitative results for multi-modal brain MRI synthesis on BraTS and IXI datasets. The numbers of the second row denote the number of missing modalities. The best performances are bolded.
BraTS IXI
1 2 3 1 2
@@
MMGAN [43] 24.71 $\pm$1.57 0.817 $\pm$0.027 27.94 $\pm$1.05 24.38 $\pm$1.74 0.806 $\pm$0.023 32.48 $\pm$1.66 24.06 $\pm$1.92 0.794 $\pm$0.031 39.37 $\pm$2.10 22.29 $\pm$1.35 0.684 $\pm$0.015 70.91 $\pm$3.64 21.13 $\pm$1.49 0.668 $\pm$0.019 93.57 $\pm$3.42
MMT [24] 25.19 $\pm$1.41 0.824 $\pm$0.017 24.53 $\pm$1.33 24.50 $\pm$1.55 0.811 $\pm$0.020 29.57 $\pm$1.82 23.92 $\pm$1.53 0.801 $\pm$0.021 39.66 $\pm$1.95 22.64 $\pm$1.49 0.698 $\pm$0.013 53.60 $\pm$3.14 21.82 $\pm$1.60 0.687 $\pm$0.019 72.24 $\pm$2.98
Hyper-GAE [53] 24.65 $\pm$1.62 0.813 $\pm$0.022 28.97 $\pm$1.57 24.42 $\pm$1.74 0.808 $\pm$0.024 33.52 $\pm$1.78 23.86 $\pm$1.88 0.788 $\pm$0.029 41.79 $\pm$2.21 22.12 $\pm$1.33 0.682 $\pm$0.017 72.62 $\pm$3.78 20.91 $\pm$1.45 0.662 $\pm$0.021 98.79 $\pm$4.05

### v2

@@ Abstract
Missing data problems, such as missing modalities in multi-modal brain MRI and missing slices in cardiac MRI, pose significant challenges in clinical practice. Existing methods rely on external guidance to supply detailed missing state for instructing generative models to synthesize missing MRIs. However, manual indicators are not always available or reliable in real-world scenarios due to the unpredictable nature of clinical environments. Moreover, these explicit masks are not informative enough to provide guidance for improving semantic consistency. In this work, we argue that generative models should infer and recognize missing states in a self-perceptive manner, enabling them to better capture subtle anatomical and pathological variations. Towards this goal, we propose CoPeDiT, a general-purpose latent diffusion model equipped with completeness perception for unified synthesis of 3D MRIs. Specifically, we incorporate dedicated pretext tasks into our tokenizer, CoPeVAE, empowering it to learn completeness-aware discriminative prompts, and design MDiT3D, a specialized diffusion transformer architecture for 3D MRI synthesis that effectively uses the learned prompts as guidance to enhance semantic consistency in 3D space. Comprehensive evaluations on three large-scale MRI datasets demonstrate that CoPeDiT significantly outperforms state-of-the-art methods, achieving superior robustness and yielding high-fidelity, structurally consistent synthesis across diverse missing patterns.
@@ 1 Introduction
Magnetic resonance imaging (MRI) provides crucial anatomical and pathological insights, particularly through multi-modal brain and volumetric cardiac scans Dickinson et al. (2013); Lustig et al. (2007); Dayarathna et al. (2024); Manna et al. (2024). However, real-world clinical MRIs frequently suffer from missing data, including absent brain modalities and missing cardiac slices, due to limited scan times, image corruption, or protocol variations Wang et al. (2025b); Paproki et al. (2024); Zhao & Shen (2025).
Figure 1: Overview and advantages of CoPeDiT. (a) We shift the paradigm from explicit mask-dependent guidance to autonomous completeness perception via prompt generation. This self-perceptive mechanism offers significant advantages: (b) quantitative plug-and-play performance gains across existing baselines; (c) substantially more discriminative latent representations compared to training without pretext tasks; and (d) enhanced semantic attention alignment between correlated modalities (e.g., T1-T1ce, T2-FLAIR) compared to traditional binary mask codes.
To address this, generative models have been developed to infer missing data from observed inputs Ibrahim et al. (2025); Ferreira et al. (2024). Existing paradigms rely on auxiliary embeddings, e.g., binary mask codes, as prior knowledge to encode missing patterns (e.g., severity, type, and position) Liu et al. (2023); Hao et al. (2024); Kim & Park (2024); Cho et al. (2024); Meng et al. (2024). Nonetheless, these hand-crafted masks only indicate missing locations, without adequately characterizing the actual incomplete state of the input. This causes three practical limitations. First, because missing patterns vary across hospitals, scanners, and acquisition settings, enumerating them with predefined masks is unrealistic in real deployments Wang et al. (2023); Rui et al. (2025). Second, the resulting condition is insensitive to modality-specific and spatially varying context, making models less robust to unseen incomplete patterns and prone to degraded generalization Lee et al. (2023); Ke et al. (2025); Wenderoth et al. (2025); Hamamci et al. (2024); Azad et al. (2025); Pan et al. (2025). Third, because binary masks carry limited semantics, they provide rigid and insufficiently informative guidance, which can weaken spatial alignment and semantic consistency during synthesis Qiu et al. (2023); Hu et al. (2023); Shin et al. (2025); Wu et al. (2025).
@@
Intuitively, generative models should be capable of inferring and detecting the incomplete state spontaneously, rather than relying on externally provided manual guidance Hu et al. (2023); Graikos et al. (2024). Motivated by this, we pose a central question: Can we empower the model with the ability to perceive missing states on its own? In light of this, we exploit an underexplored property of generative models in medical imaging, i.e., completeness perception, to enhance flexibility and generalizability under arbitrary missing MRI conditions, as illustrated in Fig. 1. Our fundamental insight is to enable the generative model to recognize the fine-grained incomplete state information in a self-perceptive manner, and to leverage this understanding as internal prompts to guide the generation process. We hypothesize that, for diffusion models, such self-guided prompts may serve as an effective alternative to manually defined masks, and potentially offer even stronger guidance signals (Fig. 1). The main reason lies in the fact that this self-perceptive strategy encourages the model to learn both global and local anatomical structures and lesion patterns at coarse and fine levels, thus enabling more semantically coherent generation of the missing MRI regions during synthesis (Figs. 11Liang et al. (2022).
@@
Driven by our motivation, we propose CoPeDiT, a 3D latent diffusion model (LDM) framework for unified 3D MRI synthesis. Here, our “unified” denotes a common formulation, training paradigm, and conditioning strategy, with task-specific adaptations for different synthesis settings. Technically, our framework builds on two core components: (i) Unlike prior approaches that require explicit missing indicators, a novel tokenizer with a completeness perception function, CoPeVAE, is proposed to autonomously assess the integrity of modalities or volumes through tailored self-supervised pretext tasks. By detecting anatomical structures and variations in lesion patterns, CoPeVAE develops a comprehensive understanding of 3D MRIs. This enables CoPeDiT to eliminate the need for manual intervention with flexible adaptability, enhancing the method’s autonomy and improving its feasibility for real-world clinical deployment with diverse missing patterns. (ii) A task-specific diffusion transformer instantiation, MDiT3D, is developed as a dependency-aligned conditioning interface for completeness-aware prompts in 3D MRI synthesis. Rather than introducing a fundamentally new DiT paradigm, MDiT3D adapts tokenization, attention, and prompt injection to the long-range, anisotropic, and irregular structural dependencies of volumetric MRIs. This design provides the proper pathway for our learned prompts to influence the generative process, allowing the prompt tokens to propagate observed modality- or slice-specific cues along anatomically meaningful dependencies during diffusion. As a result, MDiT3D enables more reliable synthesis of missing modalities or slices while better preserving structural consistency in high-dimensional 3D data. Incorporating the above two innovations, our architecture not only enables adaptive self-guidance synthesis, but also demonstrates improved structural coherence and enhanced preservation of fine-grained anatomical details. Our main contributions are summarized as follows:
@@
  • We propose a unified formulation with task-specific instantiations, dubbed CoPeDiT, for both 3D brain and cardiac missing MRI synthesis under arbitrary incomplete scenarios, without the need for explicit external indicators as guidance.
@@
  • We present MDiT3D, a tailored diffusion transformer for 3D MRI synthesis, designed as a dependency-aligned conditioning interface for our completeness-aware prompts to enable structurally meaningful prompt-guided generation.
  • Extensive experiments on three datasets demonstrate that our model surpasses state-of-the-art (SOTA) methods, achieving robustness and real-world clinical applicability.
@@ 2 Related Work
Medical Image Generation. Medical image generation has been widely studied for data augmentation Hamamci et al. (2024); Zhao et al. (2025), reconstruction Yu et al. (2025); Xiao et al. (2025); Liu et al. (2024), and image completion Rassmann et al. (2025); Liu et al. (2023); Song et al. (2026) in clinical imaging workflows. Earlier methods based on Generative Adversarial Networks (GANs) improved realism but frequently faced limitations in training stability and mode collapse, restricting their fidelity, especially when scaling to complex, high-dimensional 3D volumetric data Cao et al. (2023); Shao et al. (2025); Weng et al. (2024). To overcome these bottlenecks, diffusion-based models have emerged as a robust alternative. By offering stable training dynamics and superior mode coverage, they have achieved competitive performance in medical image synthesis and are increasingly favored in conditional settings guided by partial observations, anatomical priors, or auxiliary modalities Wang et al. (2025a); Zhao et al. (2025); Zhang et al. (2026); Guo et al. (2025); Yeganeh et al. (2025).
MRI Synthesis. Recent unified MRI synthesis methods predominantly utilize generative models, ranging from GANs Xia et al. (2021); Zhang et al. (2024; 2019a; 2019b); Yang et al. (2023); Sharma & Hamarneh (2020) and Transformers Liu et al. (2023; 2021; 2025) to advanced diffusion models Qiu et al. (2025); Yeganeh et al. (2025); Meng et al. (2024); Song et al. (2026). While these approaches successfully capture complex inter-modal dependencies to impute missing data, they inherently rely on externally provided masks to explicitly encode missing patterns in randomly incomplete scenarios Liu et al. (2023); Meng et al. (2024); Wang et al. (2023); Azad et al. (2025). This manual guidance is often rigid and lacks informative semantic details about the actual incomplete state of the input. Furthermore, because missing patterns vary widely across different clinical environments, requiring predefined masks limits practical deployment. In contrast to these mask-dependent paradigms, CoPeDiT explores the self-perceptive capability of generative models to autonomously recognize data completeness. By learning to infer missing states internally, our approach eliminates the need for manual intervention, enabling more flexible and high-fidelity MRI synthesis.
Diffusion Models. Diffusion models, particularly LDMs Rombach et al. (2022), have demonstrated remarkable capabilities across various vision tasks Ho et al. (2020); Lu et al. (2024); Ma et al. (2025); Yao et al. (2025). Recently, DiTs Peebles & Xie (2023) have emerged as powerful alternatives to traditional U-Net Ronneberger et al. (2015) backbones, achieving competitive performance in natural image synthesis. However, most existing medical image generation approaches still predominantly rely on U-Net architectures, leaving the potential of transformer-based LDMs largely underexplored in this domain Wang et al. (2025a); Nazir et al. (2025). To bridge this gap, we introduce MDiT3D, which replaces the U-Net backbone with a diffusion transformer. By incorporating task-specific architectural modifications, MDiT3D effectively captures the complex, long-range dependencies inherent in high-dimensional 3D MRIs.
3 Methodology
@@ 3.1 Notations
To unify the formulation of brain missing-modality synthesis and cardiac missing-slice synthesis, let $\mathcal{M}=\{\mathbf{x}^{i}\}_{i=1}^{m}$ denote a complete 3D MRI sample comprising $m$ elements (either modalities or slices). We partition $\mathcal{M}$ into an available subset $\mathcal{M}^{S}=\{\mathbf{x}^{s_{i}}\}_{i=1}^{s}$ and a missing subset $\mathcal{M}^{C}=\{\mathbf{x}^{c_{i}}\}_{i=1}^{c}$, where $m=s+c$. The objective is to synthesize $\mathcal{M}^{C}$ from $\mathcal{M}^{S}$. Specifically, for brain MRIs, $\mathcal{M}^{C}$ contains randomly missing modalities; for cardiac MRIs, $\mathcal{M}^{C}$ consists of consecutive missing slices. To mimic real-world clinical environments, we evaluate randomly generated incomplete cases with varying missing counts and lengths.
@@ 3.2 Stage I: Completeness Perception Tokenizer
The core idea of CoPeVAE (Fig. 2) is that detecting the completeness of high-resolution MRI data enforces the model to perceive both global anatomy and local lesion patterns, thereby producing high-quality prompts as diffusion guidance. Building upon VQGAN van den Oord et al. (2017); Esser et al. (2021), we deploy a 3D autoencoder jointly trained with self-supervised pretext tasks. Each task employs a prompt encoder, denoted as $\mathcal{F}_{1},\mathcal{F}_{2}$ and $\mathcal{F}_{3}$, to transform latent tokens (learned by the encoder $\mathcal{E}$) into low-dimensional prompt tokens. All prompt encoders contain 3D Conv layers followed by spatial average pooling. Afterwards, task-specific projection heads are applied for multi-granular classification and contrastive learning.
@@
Task 1. Missing Number/Length Detection. Equipped with global contextual awareness, the tokenizer is capable of determining how many modalities/slices are missing in the incomplete input. This task aims to enable the tokenizer to identify the severity of incompleteness by perceiving the global context of MRIs, thereby learning modality/spatial attributes in a coarse-grained manner. We formulate this task as an (m1)(m-1)-class classification task and define the loss using cross-entropy loss as follows:
$\mathcal{L}_{d}=\mathcal{L}_{\text{cls}}(\mathcal{H}_{1}(\mathcal{F}_{1}(\mathbf{z}^{s}))_{c}).$ (1)
where $\mathcal{F}_{1}$ and $\mathcal{H}_{1}$ represent the prompt encoder and the projection head for missing detection, respectively. The learned prompt tokens $\mathbf{p}^{d}=\mathcal{F}_{1}(\mathbf{z}^{s})$ are rich in information about the severity of the missing state.
Figure 3: Architecture of the MDiT3D framework. We implement two variants with task-specific alternating blocks (Spatial/Modal for MDiT3D-B and Planar/Spatial for MDiT3D-C), using dynamic feature reshaping to accurately model respective anatomical dependencies. Learned prompts are injected via adaLN as conditional guidance. PE: Positional Embeddings; RoPE: Rotary Position Embeddings Su et al. (2024); Yang et al. (2025).
@@
$\mathcal{L}_{p}=\mathcal{L}_{\text{cls}}(\mathcal{H}_{2}(\mathcal{F}_{2}(\mathbf{z}^{s}))_{I}).$ (2)
@@
Task 3. Missing Modality/Slice Assessment. Motivated by the observation that modalities or slices from the same scan share more similar anatomical and textural context than those from different scans, we adopt an inter-modal/inter-slice contrastive learning scheme Radford et al. (2021) to serve as a missing data estimator. Specifically, we take the incomplete latent tokens $\mathbf{z}^{s}$ as the anchor, the corresponding missing latent $\mathbf{z}^{c}$ from the same subject as positives, and latent tokens $\mathbf{z}^{c}_{-}$ from different subjects as negatives. The contrastive loss is defined as follows:
$\mathcal{L}_{c}=-\log\frac{\varphi\left(\mathcal{H}_{3}(\mathbf{p}^{s}),\mathcal{H}_{3}(\mathbf{p}^{c})\right)}{\varphi\left(\mathcal{H}_{3}(\mathbf{p}^{s}),\mathcal{H}_{3}(\mathbf{p}^{c})\right)+\sum\varphi\left(\mathcal{H}_{3}(\mathbf{p}^{s}),\mathcal{H}_{3}(\mathbf{p}^{c}_{-})\right)}.$ (3)
where $\mathbf{p}^{s}=\mathcal{F}_{3}(\mathbf{z}^{s})$, $\quad\mathbf{p}^{c}=\mathcal{F}_{3}(\mathbf{z}^{c})$, $\varphi(a,b)=\exp(a\cdot b/\tau)$, and $\tau$ is the temperature parameter. This contrastive learning scheme encourages the model to focus on inter-modal/slice contextual differences, improving anatomical coherence and fine-grained detail preservation. The overall loss of CoPeVAE is formulated as
$\mathcal{L}_{\text{tok}}=\mathcal{L}_{\text{rec}}+\lambda(\mathcal{L}_{d}+\mathcal{L}_{p}+\mathcal{L}_{c}).$ (4)
where $\lambda$ is the weighting coefficient, and $\mathcal{L}_{\text{rec}}$ is the reconstruction loss containing a L1 loss, vector quantization loss, adversarial loss, and perceptual loss.
Table 1: Quantitative results for multi-modal brain MRI synthesis on the BraTS dataset. The numbers in the first row denote the number of missing modalities.
1 2 3
@@
MMGAN Sharma & Hamarneh (2020) 24.71 $\pm$1.57 0.817 $\pm$0.027 27.94 489.86 24.38 $\pm$1.74 0.806 $\pm$0.023 32.48 726.93 24.06 $\pm$1.92 0.794 $\pm$0.031 39.37 898.17
MMT Liu et al. (2023) 25.19 $\pm$1.41 0.824 $\pm$0.017 24.53 527.06 24.50 $\pm$1.55 0.811 $\pm$0.020 29.57 686.52 23.92 $\pm$1.53 0.801 $\pm$0.021 39.66 841.46
Hyper-GAE Yang et al. (2023) 24.65 $\pm$1.62 0.813 $\pm$0.022 28.97 609.65 24.42 $\pm$1.74 0.808 $\pm$0.024 33.52 815.92 23.86 $\pm$1.88 0.788 $\pm$0.029 41.79 948.27